# On the Cuspy Structure of Rotating Wormhole Shadows

### v1

@@ 1 Introduction
The shadow of a compact gravitating object has emerged as a cornerstone of strong-field gravity phenomenology. Since the Event Horizon Telescope’s groundbreaking images of M87* and Sagittarius A*, the study of black hole shadows has matured into a precise and rich discipline, offering important insight into spacetime geometry of ultracompact objects [1, 2, 3, 4, 5, 6]. Yet black holes are not the only theoretically viable compact objects. More general geometries, such as naked singularities, gravastars, and wormholes, can also cast observable shadows [7, 8, 9, 10, 11]. In particular, as one of the most intriguing predictions of Einstein’s theory of gravity [12, 13], traversable wormholes, which constitute spacetime tunnels linking remote parts of the universe, are both rigorously formulated scientific models and enduring subjects of theoretical investigation. The observational signatures, particularly their shadows, provide a concrete pathway to test the black hole paradigm and to search for horizonless alternatives. Thus, the investigation of the shadow cast by traversable wormholes holds significant importance.
@@
Traversable wormholes were originally conceived by Morris and Thorne as solutions of Einstein’s gravity coupled to exotic matter [14]. To construct a singularity-free metric to describe wormhole geometry that is consistent with the Einstein field equations, the matter contained in Morris-Thorne wormholes must possess a stress-energy tensor that violates the null energy condition. Subsequently, rotating axially symmetric solutions describing traversable wormholes were also obtained [15, 16]. The apparent violation of the energy conditions may be either mitigated or circumvented depending on the theoretical context [17, 18, 19, 20, 21, 22], which establishes the traversable wormhole as a robust and conceptually rich model. In this work, we will study the shadow of a general class of rotating traversable wormholes in the Teo class [15].
@@
A further reason for studying the shadow cast by rotating wormholes is due to the increasing attention on the cuspy shadow boundaries [23, 24, 25, 26, 27, 28]. The cusp is not numerical noise, but a distinctive structure dictated by the photon orbit configuration in a strong gravitational system. The recent study [23] reveals that the formation of cusps in black hole shadows constitutes a topological phase transition and a gravitational equal-area law. Moreover, the critical point of cusp formation has universal exponents, placing the phenomenon in the mean-field universality class. The study implies that the emergence of cuspy structures is endowed with deep physical meaning, rendering them theoretically significant. The shadow of the rotating traversable wormhole has also been reported to exhibit a cuspy edge [10, 11]. Thus, we are naturally eager to comprehensively understand the cuspy structure of the wormhole shadow. However, early investigations of rotating wormhole shadows do not provide a careful enough study on the formation of the cuspy edge [9, 10, 11, 29, 30, 31]. Especially, most of these studies have adopted the simplest possible redshift function, typically $N=\exp(−r_{0}/r)$ or $N=\exp(−r_{0}^{2}/r^{2})$. As will be illustrated in this work, those models are too simple to properly capture the process of cusp formation. In principle, the redshift function $N(r)$ can be chosen more freely, like
@@
As will be demonstrated in this paper, the shadow of rotating traversable wormholes exhibits a surprisingly rich structure governed by the redshift parameter $\lambda$ and spin parameter $a$. More explicitly, we uncover a universal critical value $\lambda_{c}=(\sqrt{5}-1)/4$ independent of spin and throat radius, below which the shadow boundary is smooth and above which a sharp cusp emerges at the intersection of the outer circular orbits and the throat critical orbits. The formation of cusps can not be seen by the models with $N=\exp(−r_{0}/r)$ or $N=\exp(−r_{0}^{2}/r^{2})$ [9, 10, 11]. By scanning the $(\lambda,a)$ parameter space, a comprehensive phase diagram is constructed, revealing four distinct shadow morphologies: smooth, cuspy, ears touching, and throat drowning. As we vary the parameters, the shadow can exhibit interesting re-entrant behavior where the throat orbit detaches and later reconnects. The cusp boundary enables a better understanding of the photon orbits in the wormhole background, and the special morphology may serve as a diagnostic signature for distinguishing traversable wormholes from black holes and other compact objects.
The paper is organized as follows. Sec. 2 introduces the rotating traversable wormhole metric and specifies the class of solutions studied. We derive the null geodesic equations, the conditions for critical orbits, and the mapping to the observer’s celestial coordinates in Sec. 3. Then in Sec. 4, we investigate the shadow morphology with different spin and redshift parameters, revealing a universal critical redshift parameter and a phase diagram comprising four distinct phases. Sec. 5 concludes the whole paper and discusses future directions. Appendix A provides a parametric description of the wormhole throat orbit.
@@ 3 Shadow of the wormholes
    We have an explicit expression for the Bardeen impact parameter $(\alpha,\beta)$ for the throat orbits as displayed in (40), however, it may be useful to have a parametric expression for the throat orbit just like (37) as well. We discuss the parametric expression for the throat orbit in Appendix A by introducing an effective radius $\bar{r}\in[0,r0]$.
@@ 4.1 The development of the cusp with different $\lambda$
For $\lambda<\lambda_{c}$, for example $\lambda=0.2$, the light from the unstable circular orbits and the wormhole throat form a smooth curve, as shown in Fig. 2a. We can prove that the outer circular orbit and throat boundary approach each other with the same slope. We define the slope of the shadow boundary, following [23], as
@@
So the slope of the outer circular orbits when approaching the throat orbit can be obtained from (37)
@@
From (38), the slope of the throat orbit when it approaches the outer circular orbits is
@@
The critical point appears at $\lambda=\lambda_{c}$, as shown in Fig. 2b. For values exceeding this threshold ($\lambda>\lambda_{c}$), a cuspy structure develops and becomes increasingly distinct. We depict the shadow with $\lambda=0.8$ in Fig. 2c, where the “swallowtail” behavior of the outer unstable circular orbits is explicitly demonstrated. It is worth noting that the swallowtail structure itself, though present in the diagram, is not directly visible in the observational shadow. There can be light coming from the triangle region enclosed by the red and blue curves shown in Fig. 3. What we can observe is a cusp of the shadow located at $(\alpha,\beta)=(-2.79,2.41)$, for the case with $a=0.07$ and $\lambda=0.8$. The outer unstable circular orbit and throat orbit meet at the purple dot $(\alpha,\beta)=(-2.48,3.06)$, as demonstrated in Fig. 3. Additionally, the slopes of the two curves are the same, as discussed in equation (46). The cusp where the two curves intersect is illustrated by the red dot in the figure. The slopes of the two curves are different at the cusp point, and we have a slope discontinuity, hence the term “cusp”.
Figure 3: The cusp and swallowtail for the unstable circular orbits, with $a=0.07$ and $\lambda=0.8$. The outer unstable circular orbit (blue) and throat orbit (red) meet at the purple dot, with the same slope $\mathcal{F}_{\text{out}}=\mathcal{F}_{\text{throat}}$. The cusp is illustrated by the red dot.
For larger redshift parameter $\lambda$, the “ears” of the swallowtail start to touch each other at a critical value $\lambda_{\text{touch}}\approx 0.99$, as shown in Fig. 2d. Although this is not an observable behavior, the phenomenon is interesting to study. At $\lambda_{\text{touch}}$, the blue curve representing the outer unstable circular orbit, seals off into a quasi-circular shape. As displayed in Fig. 2d, as we raise the photon orbit radius $r_{\text{ph}}$ from $r_{0}=1$, $\beta$ decreases to zero and then starts to increase afterward. For $\lambda>\lambda_{\text{touch}}$, the blue curve breaks into two disjoint segments: a quasi-circular curve and a curve connecting the throat orbit, which is plotted in Fig. LABEL:63.
As we keep increasing $\lambda$, the throat orbit represented by the red curve would decouple from the blue quasi-circular curve, such that the shadow is solely enveloped by the outer unstable orbit, as displayed in Fig. 4a. The wormhole throat recedes from view and does not play any role in the shadow. We can call this situation “throat drowning” and denote the critical value of $\lambda$ as $\lambda_{\text{drown}}$. For even larger $\lambda>\lambda_{\text{drown}}$, for instance $\lambda=1.7$ shown in Fig. LABEL:42, the triangle region enclosed by the blue and red curves separates from the wormhole shadow. One may expect the region related to the wormhole throat to be separated forever. This expectation, however, is not borne out, and we can see from Fig. 4c that the wormhole throat orbit reconnects with the shadow again at $\lambda=5.41$. For $\lambda>5.41$, the wormhole contributes to the shadow boundary again, and the shadow becomes the common envelope of the two curves again, for large $\lambda$. Note that for large $\lambda$, the throat orbit behaves like a vertical line with fixed $\alpha$ as shown in Fig. 4d. This tendency becomes increasingly pronounced as $\lambda$ grows further. The throat-drowning situation, however, is not common for all rotating wormholes. As will be demonstrated subsequently, the throat orbit submerges only for sufficiently small spin parameters. For spin parameter $a>0.08$, the wormhole throat would invariably contribute to the shadow boundary.
@@
Figure 4: The rotating traversable shadow with $\lambda$ (with $a=0.07$). For $\lambda\geq\lambda_{\text{drown}}$, with increasing $\lambda$, the wormhole throat orbit first detaches from the shadow, and then becomes tangent with the shadow at $\lambda=5.41$. For even larger $\lambda$, the shadow boundary comprises the throat orbits and the outer unstable circular orbits.
@@ 4.2 Phase diagram of the shadow boundary
The diagram representing different phases of the wormhole shadow boundary is illustrated in Fig. 5. As can be seen from the figure, for $\lambda<\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$, no matter what the spin parameter $a$ is, the shadow boundary is smooth, and the shadow exhibits no cusp structure. This is consistent with previous studies where the model with $N=\exp(-r_{0}/r)$ is considered [10, 11]. As $\lambda$ continues to increase, the cusp starts to form at the critical value $\lambda_{c}$. For larger $\lambda$, there are three different phases, which are cuspy shadow, ears touching, and throat drowning, respectively. The blue dotted line in Fig. 5 illustrates different phases with fixed $\lambda=2$, and the corresponding morphology of the shadow boundary is displayed in Fig. 6. The red dotted line represents the situations discussed in the previous subsection with $a=0.07$. As discussed before, for fixed $a=0.07$, the cusps start to appear at $\lambda=\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$ as we increase $\lambda$. Keep increasing $\lambda$, the ears of the shadow boundary start to get connected at $\lambda=\lambda_{\text{touch}}\approx 0.99$. For $1.38<\lambda<5.54$, the throat can not be seen by the shadow observation, and we are in the throat-drowning region. The shadow is determined solely by the outer unstable orbit and is smooth in this region. For $\lambda>5.54$, the wormhole throat orbit gets reconnected to the shadow, and the shadow becomes enclosed by the two orbits with cusps, as shown in Fig. 4d.
@@
  • For a rotating wormhole with a large spin parameter, there are cuspy structures for the wormhole shadow, and the outer unstable circular orbit exhibits swallowtail behavior as shown in Fig. 6a and Fig. 3. In this case, the shadow is the shaded region enclosed by the wormhole throat orbit and the outer unstable orbit.
  • In the ears touching region, the shadow boundary is a composite contour that consists of two distinct blue and red segments, as shown in Fig. LABEL:63. However, as we continue decreasing the spin parameter, after a threshold of $a$, the wormhole throat orbit recedes from the shadow. For the $\lambda=2$ case, the critical value of the spin parameter is $a\approx 0.078$, which is displayed in Fig. LABEL:64. We call the region where the throat orbit detaches from the shadow the throat-drowning region, which is represented by the blue area in Fig. 5. The line between the ears-touching and the throat-drowning regions demonstrates the situation when the red curve starts to detach from the shadow shown in Fig. LABEL:64. There would not be a throat-drowning situation for high-spinning wormholes, regardless of the value of $\lambda$. The line between the yellow and blue regions has a maximum at $(\lambda,a)\approx(2.68,0.08)$.
  • In the throat-drowning region, the wormhole throat orbit never contributes to the shadow. Moreover, for very small $a$, the wormhole throat orbit may shrink to a single point as displayed in Fig. 7. For the case $\lambda=2$, the critical value for wormhole throat orbit shrinking to a point is $a\approx 0.033$. It is worth emphasizing the difference between situations illustrated in Fig. LABEL:64 and in Fig. 7. Both cases are in the throat-drowning region, meaning the wormhole throat orbit would not contribute to the wormhole shadow. For the case with $0.033\leq a\leq 0.078$, the wormhole throat orbit detaches from the shadow, thus it can not be observed from the shadow. However, the throat orbit is still there, and one may receive light from the wormhole throat. For $a<0.033$, the wormhole throat orbit completely disappears, and no light signal from the throat can be received.
@@
It is important to describe the overall morphology and global deformation of the shadow as the spin parameter $a$ and the redshift parameter $\lambda$ vary. The transition from a smooth to a cuspy shadow boundary, and indeed the very formation of the cusp itself, only emerges upon varying the redshift parameter $\lambda$. As shown in Fig. 2, the size of the shadow grows as we increase $\lambda$. The spin parameter does not have much influence on the size of the shadow, as can be seen from Fig. 6. The primary role of the spin parameter is to modulate the horizontal asymmetry of the shadow about the vertical axis. In Fig. 6, for small $a$ like figures LABEL:63 and LABEL:64, the shadow exhibits near-perfect left-right symmetry about the y-axis. While for relatively larger $a$, the wormhole shadow exhibits a pronounced displacement in the positive $\alpha$ direction, see Fig. 6a for example. We have discussed that for large $\lambda$, the wormhole throat orbit behaves like a vertical line. Moreover, the size of the swallowtail boundary also increases for large values of $\lambda$. This contrast is further illustrated in Fig. 8, where two representative cases are shown. For $a=0.95$ and $\lambda\approx 15.33$, the size of the swallowtail is huge, and the shadow exhibits a semicircular disk morphology. For parameters $a=0.05$ and $\lambda\approx 0.67$, the swallowtail can barely be seen, and the shadow is symmetrically distributed. The close relation between the shadow morphology and the two parameters $\lambda$ and $a$, provides a clear observational signature that may help distinguish the two parameters from future high-resolution images.
@@
Figure 7: The wormhole shadow with $\lambda=2$ and $a\approx 0.033$ . For this critical value of spin parameter, the wormhole throat orbit shrinks to a single point. The throat orbit completely disappears for even smaller $a$.
@@
For small $a<0.08$, we have seen a re-entrant behavior of the throat orbit, detachment followed by reattachment, to the shadow. Notably, for relatively large $\lambda$, the throat orbit reconnects with the shadow and behaves almost as a vertical line in the $(\alpha,\beta)$ plane, while the outer orbit continues to expand. This phenomenon can be understood as a non-trivial competition between the redshift effect (which enlarges the shadow) and the frame-dragging effect (which distorts its shape).

### v2

@@ 1 Introduction
The shadow of a compact gravitating object has emerged as a cornerstone of strong-field gravity phenomenology. Since the Event Horizon Telescope’s groundbreaking images of M87* and Sagittarius A*, the study of black hole shadows has matured into a precise and rich discipline, offering important insight into spacetime geometry of ultracompact objects [1, 2, 3, 4, 5, 6]. Yet black holes are not the only theoretically viable compact objects. More general geometries, such as naked singularities, gravastars, and wormholes, can also cast observable shadows [7, 8, 9, 10, 11]. In particular, as one of the most intriguing predictions of Einstein’s theory of gravity [12, 13], traversable wormholes, which constitute spacetime tunnels linking remote parts of the universe, are both rigorously formulated scientific models and enduring subjects of theoretical investigation. The observational signatures, particularly their shadows, provide a concrete pathway to test the black hole paradigm and to search for horizonless alternatives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Thus, the investigation of the shadow cast by traversable wormholes holds significant importance.
@@
Traversable wormholes were originally conceived by Morris and Thorne as solutions of Einstein’s gravity coupled to exotic matter [27]. To construct a singularity-free metric to describe wormhole geometry that is consistent with the Einstein field equations, the matter contained in Morris-Thorne wormholes must possess a stress-energy tensor that violates the null energy condition. Subsequently, rotating axially symmetric solutions describing traversable wormholes were also obtained [28, 29]. The apparent violation of the energy conditions may be either mitigated or circumvented depending on the theoretical context [30, 31, 32, 33, 34, 35, 36, 37, 38], which establishes the traversable wormhole as a robust and conceptually rich model. In this work, we will study the shadow of a general class of rotating traversable wormholes in the Teo class [28].
@@
A further reason for studying the shadow cast by rotating wormholes is due to the increasing attention on the cuspy shadow boundaries [39, 40, 41, 42, 43, 44]. The cusp is not numerical noise, but a distinctive structure dictated by the photon orbits in a strong gravitational system. The recent study [39] reveals that the formation of cusps in black hole shadows constitutes a topological phase transition and a gravitational equal-area law. Moreover, the critical point of cusp formation has universal exponents, placing the phenomenon in the mean-field universality class. The study implies that the emergence of cuspy structures is endowed with deep physical meaning, rendering them theoretically significant. The shadow of the rotating traversable wormhole has also been reported to exhibit a cuspy edge [10, 11]. Thus, we are naturally eager to comprehensively understand the cuspy structure of the wormhole shadow. However, early investigations of rotating wormhole shadows do not provide a careful enough study on the formation of the cuspy edge [9, 10, 11, 45, 46, 47, 48]. Especially, most of these studies have adopted the simplest possible redshift function, typically $N=\exp(−r_{0}/r)$ or $N=\exp(−r_{0}^{2}/r^{2})$. As will be illustrated in this work, those models are too simple to properly capture the process of cusp formation. In principle, the redshift function $N(r)$ can be chosen more freely, like
@@
As will be demonstrated in this paper, the shadow of rotating traversable wormholes exhibits a surprisingly rich structure governed by the redshift parameter $\lambda$ and spin parameter $a$. More explicitly, we uncover a universal critical value $\lambda_{c}=(\sqrt{5}-1)/4$ independent of spin and throat radius, below which the shadow boundary is smooth and above which a sharp cusp emerges at the intersection of the outer circular orbits and the throat critical orbits. The formation of cusps can not be seen by the models with $N=\exp(−r_{0}/r)$ or $N=\exp(−r_{0}^{2}/r^{2})$ [9, 10, 11]. By scanning the $(\lambda,a)$ parameter space, a comprehensive phase diagram is constructed, revealing four distinct shadow morphologies: smooth, cuspy, ears touching, and throat drowning. As we vary the parameters, the shadow can exhibit interesting re-entrant behavior where the curve of the throat orbits detaches and later reconnects. The cusp boundary enables a better understanding of the photon orbits in the wormhole background, and the special morphology may serve as a diagnostic signature for distinguishing traversable wormholes from black holes and other compact objects.
The paper is organized as follows. Sec. 2 introduces the rotating traversable wormhole metric and specifies the class of solutions studied. We derive the null geodesic equations, the conditions for critical orbits, and the mapping to the observer’s celestial coordinates in Sec. 3. Then in Sec. 4, we investigate the shadow morphology with different spin and redshift parameters, revealing a universal critical redshift parameter and a phase diagram comprising four distinct phases. Sec. 5 concludes the whole paper and discusses future directions. Appendix A provides a parametric description of the wormhole throat orbits.
@@ 3 Shadow of the wormholes
    We have an explicit expression for the Bardeen impact parameter $(\alpha,\beta)$ for the throat orbits as displayed in (40), however, it may be useful to have a parametric expression for the throat orbits just like (37) as well. We discuss the parametric expression for the throat orbits in Appendix A by introducing an effective radius $\bar{r}\in[0,r0]$.
@@ 4.1 The development of the cusp with different $\lambda$
For $\lambda<\lambda_{c}$, for example $\lambda=0.2$, the lights from the unstable circular orbits and the wormhole throat form a smooth curve, as shown in Fig. 2a. We can prove that the outer circular orbits and throat boundary approach each other with the same slope. We define the slope of the shadow boundary, following [39], as
@@
So the slope of the outer circular orbits when approaching the throat orbits can be obtained from (37)
@@
From (38), the slope of the throat orbits when it approaches the outer circular orbits is
The critical point appears at $\lambda=\lambda_{c}$, as shown in Fig. 2b. For values exceeding this threshold ($\lambda>\lambda_{c}$), a cuspy structure develops and becomes increasingly distinct. We depict the shadow with $\lambda=0.8$ in Fig. 2c, where the “swallowtail” behavior of the outer unstable circular orbits is explicitly demonstrated. It is worth noting that the swallowtail structure itself, though present in the diagram, is not directly visible in the observational shadow. There can be light coming from the triangle region enclosed by the red and blue curves shown in Fig. 3. What we can observe is a cusp of the shadow located at $(\alpha,\beta)=(-2.79,2.41)$, for the case with $a=0.07$ and $\lambda=0.8$. The outer unstable circular orbits and throat orbits meet at the purple dot $(\alpha,\beta)=(-2.48,3.06)$, as demonstrated in Fig. 3. Additionally, the slopes of the two curves are the same, as discussed in equation (46). The cusp where the two curves intersect is illustrated by the red dot in the figure. The slopes of the two curves are different at the cusp point, and we have a slope discontinuity, hence the term “cusp”.
Figure 3: The cusp and swallowtail for the unstable circular orbits, with $a=0.07$ and $\lambda=0.8$. The outer unstable circular orbits (blue) and throat orbits (red) meet at the purple dot, with the same slope $\mathcal{F}_{\text{out}}=\mathcal{F}_{\text{throat}}$. The cusp is illustrated by the red dot.
For larger redshift parameter $\lambda$, the “ears” of the swallowtail start to touch each other at a critical value $\lambda_{\text{touch}}\approx 0.99$, as shown in Fig. 2d. Although this is not an observable behavior, the phenomenon is interesting to study. At $\lambda_{\text{touch}}$, the blue curve representing the outer unstable circular orbits, seals off into a quasi-circular shape. As displayed in Fig. 2d, as we raise the photon orbit radius $r_{\text{ph}}$ from $r_{0}=1$, $\beta$ decreases to zero and then starts to increase afterward. For $\lambda>\lambda_{\text{touch}}$, the blue curve breaks into two disjoint segments: a quasi-circular curve and a curve connecting the throat orbits, which is plotted in Fig. LABEL:63.
@@
Figure 4: The rotating traversable shadow with $\lambda$ (with $a=0.07$). For $\lambda\geq\lambda_{\text{drown}}$, with increasing $\lambda$, the curve of the wormhole throat orbits first detaches from the shadow, and then becomes tangent with the shadow at $\lambda=5.41$. For even larger $\lambda$, the shadow boundary comprises the throat orbits and the outer unstable circular orbits.
As we keep increasing $\lambda$, the throat orbits represented by the red curve would decouple from the blue quasi-circular curve, such that the shadow is solely enveloped by the outer unstable orbits, as displayed in Fig. 4a. The wormhole throat recedes from view and does not play any role in the shadow. We can call this situation “throat drowning” and denote the critical value of $\lambda$ as $\lambda_{\text{drown}}$. For even larger $\lambda>\lambda_{\text{drown}}$, for instance $\lambda=1.7$ shown in Fig. LABEL:42, the triangle region enclosed by the blue and red curves separates from the wormhole shadow. One may expect the region related to the wormhole throat to be separated forever. This expectation, however, is not borne out, and we can see from Fig. 4c that the wormhole throat orbits reconnect with the shadow again at $\lambda=5.41$. For $\lambda>5.41$, the wormhole contributes to the shadow boundary again, and the shadow becomes the common envelope of the two curves again, for large $\lambda$. Note that for large $\lambda$, the throat orbits behave like a vertical line with fixed $\alpha$ as shown in Fig. 4d. This tendency becomes increasingly pronounced as $\lambda$ grows further. The throat-drowning situation, however, is not common for all rotating wormholes. As will be demonstrated subsequently, the throat orbits submerge only for sufficiently small spin parameters. For spin parameter $a>0.08$, the wormhole throat would invariably contribute to the shadow boundary.
@@ 4.2 Phase diagram of the shadow boundary
The diagram representing different phases of the wormhole shadow boundary is illustrated in Fig. 5. As can be seen from the figure, for $\lambda<\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$, no matter what the spin parameter $a$ is, the shadow boundary is smooth, and the shadow exhibits no cusp structure. This is consistent with previous studies where the model with $N=\exp(-r_{0}/r)$ is considered [10, 11]. As $\lambda$ continues to increase, the cusp starts to form at the critical value $\lambda_{c}$. For larger $\lambda$, there are three different phases, which are cuspy shadow, ears touching, and throat drowning, respectively. The blue dotted line in Fig. 5 illustrates different phases with fixed $\lambda=2$, and the corresponding morphology of the shadow boundary is displayed in Fig. 6. The red dotted line represents the situations discussed in the previous subsection with $a=0.07$. As discussed before, for fixed $a=0.07$, the cusps start to appear at $\lambda=\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$ as we increase $\lambda$. Keep increasing $\lambda$, the ears of the shadow boundary start to get connected at $\lambda=\lambda_{\text{touch}}\approx 0.99$. For $1.38<\lambda<5.54$, the throat can not be seen by the shadow observation, and we are in the throat-drowning region. The shadow is determined solely by the outer unstable orbits and is smooth in this region. For $\lambda>5.54$, the wormhole throat orbits get reconnected to the shadow, and the shadow becomes enclosed by the two families of orbits with cusps, as shown in Fig. 4d.
  • For a rotating wormhole with a large spin parameter, there are cuspy structures for the wormhole shadow, and the outer unstable circular orbits exhibit swallowtail behavior as shown in Fig. 6a and Fig. 3. In this case, the shadow is the shaded region enclosed by the wormhole throat orbits and the outer unstable orbits.
  • In the ears touching region, the shadow boundary is a composite contour that consists of two distinct blue and red segments, as shown in Fig. LABEL:63. However, as we continue decreasing the spin parameter, after a threshold of $a$, the wormhole throat orbits recede from the shadow. For the $\lambda=2$ case, the critical value of the spin parameter is $a\approx 0.078$, which is displayed in Fig. LABEL:64. We call the region where the throat orbits detach from the shadow the throat-drowning region, which is represented by the blue area in Fig. 5. The line between the ears-touching and the throat-drowning regions demonstrates the situation when the red curve starts to detach from the shadow shown in Fig. LABEL:64. There would not be a throat-drowning situation for high-spinning wormholes, regardless of the value of $\lambda$. The line between the yellow and blue regions has a maximum at $(\lambda,a)\approx(2.68,0.08)$.
  • In the throat-drowning region, the wormhole throat orbits never contribute to the shadow. Moreover, for very small $a$, the wormhole throat orbits may shrink to a single point as displayed in Fig. 7. For the case $\lambda=2$, the critical value for wormhole throat orbits shrinking to a point is $a\approx 0.033$. It is worth emphasizing the difference between situations illustrated in Fig. LABEL:64 and in Fig. 7. Both cases are in the throat-drowning region, meaning the wormhole throat orbits would not contribute to the wormhole shadow. For the case with $0.033\leq a\leq 0.078$, the wormhole throat orbits detach from the shadow, thus they can not be observed from the shadow. However, the throat orbits are still there, and one may receive light from the wormhole throat. For $a<0.033$, the wormhole throat orbits completely disappear, and no light signal from the throat can be received.
@@
It is important to describe the overall morphology and global deformation of the shadow as the spin parameter $a$ and the redshift parameter $\lambda$ vary. The transition from a smooth to a cuspy shadow boundary, and indeed the very formation of the cusp itself, only emerges upon varying the redshift parameter $\lambda$. As shown in Fig. 2, the size of the shadow grows as we increase $\lambda$. The spin parameter does not have much influence on the size of the shadow, as can be seen from Fig. 6. The primary role of the spin parameter is to modulate the horizontal asymmetry of the shadow about the vertical axis. In Fig. 6, for small $a$ like figures LABEL:63 and LABEL:64, the shadow exhibits near-perfect left-right symmetry about the y-axis. While for relatively larger $a$, the wormhole shadow exhibits a pronounced displacement in the positive $\alpha$ direction, see Fig. 6a for example. We have discussed that for large $\lambda$, the wormhole throat orbits behave like a vertical line. Moreover, the size of the swallowtail boundary also increases for large values of $\lambda$. This contrast is further illustrated in Fig. 8, where two representative cases are shown. For $a=0.95$ and $\lambda\approx 15.33$, the size of the swallowtail is huge, and the shadow exhibits a semicircular disk morphology. For parameters $a=0.05$ and $\lambda\approx 0.67$, the swallowtail can barely be seen, and the shadow is symmetrically distributed. The close relation between the shadow morphology and the two parameters $\lambda$ and $a$, provides a clear observational signature that may help distinguish the two parameters from future high-resolution images.
@@
Figure 7: The wormhole shadow with $\lambda=2$ and $a\approx 0.033$ . For this critical value of spin parameter, the wormhole throat orbits shrink to a single point. The throat orbits completely disappear for even smaller $a$.
@@
For small $a<0.08$, we have seen a re-entrant behavior of the throat orbits, detachment followed by reattachment, to the shadow. Notably, for relatively large $\lambda$, the throat orbits reconnect with the shadow and behave almost as a vertical line in the $(\alpha,\beta)$ plane, while the outer orbits continues to expand. This phenomenon can be understood as a non-trivial competition between the redshift effect (which enlarges the shadow) and the frame-dragging effect (which distorts its shape).